OpenStudy (anonymous):

simplify fraction (2yx^-1)^3 over y^-1

OpenStudy (anonymous):

=y(2yx^-1)^3=y*2^3y^3(x^-1)^3=y*8y^3x^-3=8x^-3y^4=(8y^4)/x^3